OpenStudy (anonymous):

How do you graph a system of constraints, find all vertices, and evaluate the objective function at each vertex to find the maximum or minimum value.?

OpenStudy (anonymous):

just like you described. the constraints (if linear) can be easily graphed. For each pair of lines, find the intersection (that will be a vertex). check the functional value at each vertex and compare to find the min or max.

OpenStudy (anonymous):

can youmshow me how like an example?

OpenStudy (anonymous):

do you want me to give you an equation?

OpenStudy (anonymous):

suppose you have \[y \ge3, x \le5, x+y=2\] as constraint and \[f \left( x,y \right)=x+2y\] as the function you want to maximize.

OpenStudy (anonymous):

sorry, let \[y \le3, x \le5, x+y \le6\] be your constraints

OpenStudy (anonymous):

 also, \[y \ge0, x \ge0.\]

OpenStudy (anonymous):

you only need to find the points of intersection and evaluate f(x,y) at those points and compare to find the max or min as required.

OpenStudy (anonymous):

this is the equation

OpenStudy (anonymous):

that kind of confused me a bit cause these equations are a little simpler

OpenStudy (anonymous):

so graph each as if it were = instead of \[\le \text{ or } \ge\]

OpenStudy (anonymous):

sorry.

OpenStudy (anonymous):

ok and its ok

OpenStudy (anonymous):

whats next?

OpenStudy (anonymous):